Project

General

Profile

Face to Face Meeting 2015-1-15

Meeting 2015-1-15

Date: 15th January 2015
Time: 14:00 UK time
Attendance: Melissa, Craig, Paul k, Chris Heidt, David A, Paulo, Pierrick, Ed O, Alan B, Geoff B, Ray, Chris R (late)
Apologies: Adam

Agenda and Minutes

Actions from Last Meeting

  • GB to add all documentation to mice Mine - GB - ALL in Tracker Documents
  • MU to look in what to release as MICE notes and how - MU - ONGOING WAITING ON DOCS
  • Kalman doesnt work on straights. - Chris and Adam looking to fix - ONGOING
  • Not yet able to test against multiple tracks per trigger with MC - Durga - ONGOING
  • Real data unpacking broken - DA - it breaks between v7.5 and 7.6 as unpacking sw not versioned its hard to pin down. He thinks he has traced the fault. Stick with v7.3 in control room with a patch for OM plots (Melissa and Rhys). Single station test data can be used to help discover how to fix this.
  • Face to face agenda - MU - DONE
  • Make mock data challenge wiki page for prep discussion. - MU - DONE
  • Check with R. Bayes that this does not affect things in main CDB geometry - C. Heidt - Polished up geometry from C Hunt need to get files from Geoff. - New action GB and MU - Geoff and Melissa meet with Jason tomorrow. - DONE
  • Ensure central fibre is correctly defined in geometry and with Tracker group and correct if necessary - C Heidt CARRIES OVER
  • Issues found with NIM paper description of external waveguide end to end mapping check with Geoff and SW - MU AT MEETING tomorrow with Geoff - ONGOING

Status of commissioning work and ongoing plan
  • Cryostats in trench
  • leak checking
  • He line purge
  • valve change - that was just supply line on two so thats now fixed.
  • waveguides
  • DAQ commissioning

Downstream cryos 3 and 4 are at 5.9 and 6.
Pins put in for waveguides and glue to fix the broken boots on the waveguides. Geoff needs details of what not to use.
Use calibrated torque wrenches in black boxes.
US cryo2 is as gone from room temp to 9 degrees in 20hours.
US cryo1 on its second cycle.
US cryos and services are in trench.
Craig thinks there is a small leak on cryo 3 at 0.1 degree rise per week.
Leaks found at fitting to change them to swageloks they are being ordered.
Missing o rings replaced.
spare vme controller order requested.
New Oxfords not yet arrived.

Waveguides are very tight and there is a concern that they will not fit through the pry. We need to make our spares in the longer length. But for Step iv we don't have time to make new connectors or manufacture new D0 ends. Things are very tight for step IV. Hook up will happen over the next few days. Kevin may come up to help. will definitely need to reposition patch panel connectors to get them to fit.

LED system:
Need a benchmark of how much light we need before Ed can update it. So we will use single LED hooked up to pulse generator, run that find out how much more light to adapt the pulser for and then that work can be done. But it depends on quality of readout. Need temp and noise ok before we do LEDs. Need 1.5 PE from LED (anywhere from 1-2).

LVDS cables may have some ground loops that cause high noise. Didn't see the high noise in single station or tracker lab tests. affecting analogues signal but not digital.

Trigger Ed has updated Yordans trigger so we should be good to go. LED trigger will go into Geneva 1 trigger.

US cryos are already cabelled up, DS is being worked on but it seems that there is an intermittent short where certain boards prefer particular slots.

Paul arrives tomorrow so fixing a lot of this stuff like what crates to put where grounds etc he will help with that and a lot of work will happen over the weekend.

Online monitoring and reconstruction will be done at a minimal levels just hits and event displays and ADC/TDC from Rhys. David will send code to Rhys but not sure how much it will help.

Melissa to help with WGS and ask Kevin and Celeste as well.

Plan for mock data challenge. see Mock Data Challenge Plan

David

What and when are there opportunities to do before step IV and what can we do to be ready for them?

  • 7th March ISIS run? (validation of double dip run)

Its just one day for us (the rest is for engineers) and its a weekend and they don't want us to run but we are trying to get to be able to use it but its a bit unknown.

Goals:
  • To synchronise with the ISIS RF which was done at the single station test in two days (but we can try to complete in 1) and we will need a day with LEDs in advance. One station in one tracker at least necessary for this but one in each trackers is best.
  • And to ensure that everything not completed as part of the mock data challenge is completed.

Alignment will not be able to be done before this and therefore it is no a goal for this run.

Speeding up the DAQ

  • Paul arrives tomorrow to inspect the firmware as it currently is and check the discriminator output matches the ADC output as that is what we will limit ourselves to if we remove analogue data from DAQ. He will return to update the firmware in this regard.

We have to set 32 discriminators at once and they are all about the same gain and pedestal, Set discriminator rate to 1%.

Can we convince ourselves that just running with discriminators is safe in which case we wont lose any data and can run as fast as we want. We need to have a level two trigger on TOF data (timing cut etc) to have an analysable muon selection.

Need to have a new working group of all parties to discuss

Step IV Readiness, to do list, timeline etc.

Discuss at Mock data challenge post mortem meeting on 28th January 3pm.

Risks
h4. Hardware

  • Is it wise to get the spare VME controller from Strathclyde (they are happy for this) just in case for mock data run? - No we have ordered a spare and it has been approved.
  • Daresbury electronics: vacuum, tracker interlocks
  • Reforming external waveguides (could lead to breakage)
  • Northside PRY (not us, but knock effect to us)
  • Power supply (spare arrives mid Jan) - MITIGATED

8.2 Software - SEE ACTIONS

  • Not yet able to test against multiple tracks per trigger with MC -Action for Durga
  • Real data unpacking broken DA

Tracker Online Emergency Procedure

  • Provide documentation for ISIS and eventually shifters in what to do in an emergency situation for the Tracker when running
    Documentation to Include:
    - schematics of the system
    - a list of associated checklists at power on, and who works through them.
    - a list of instructions to shifters on how to operate and monitor the system
    - emergency shutdown instructions
    - power-up instructions
    - trouble-shooting : what to watch for
    - a list of regular maintenance and safety checks
  • If the system shuts down do not restart Tracker expert only.
  • Power outage greatest risk, cryostat will stay cold lid will be cold and moisture will get on electronics. - If this happens run lid heater boxes of UPS.

SCHEMATIC OF SYSTEM
From NIM paper

ASSOCIATED CHECKLIST AT POWER ON AND RESPONSIBILITY
Tracker expert

LIST FOR SHIFTERS OF HOW TO OPERATE AND MONITOR TRACKER SYSTEM
  • ongoing consider post Mock Data Challenge
Emergency shutdown Procedure
  • Turn off veners
  • Isolate cryostat vacuum
  • Turn off pumps
  • Turn off compressors.
IDENTIFY ITEMS THAT NEED REGULAR MAINTENANCE AND SAFETY CHECKS
  • Cryostats
  • Cold heads
  • Helium bottles

AoB

Actions

  • Send waveguide details of broken WGs to Geoff from QA - MU
  • Spare waveguides should be longer than existing long ones says Geoff
  • Melissa to help with WGS and ask Kevin and Celeste as well. _ MU
  • Prepare a tag out lock out procedure to stop people turning things off that we don't want them to (He etc.).
  • Add the above to the controls and monitoring alarms so that if the above happens we know - Pierrick
  • Add to John Govans rota for daily walk round to test our setup. - Pierrick
  • Email Yordan Chris R, Victoria and Tracker group to set up a working group for level 2 trigger - MU.
  • Plan to set up working group for level 2 trigger as a result of above.
  • make a request for UPS on heater box lids - Pierrick
  • write emergency shutdown procedure - Alan B. and David A.

Phone Details

  • The four digit company id: 4673 (callers from outside the UK)
  • The pin number: 419357

Phone numbers:

UK (Imperial College --preferred) 0845 545 4460 English
Austria 0820 4000 1503 English
0820 4000 15 02 German
Belgium 070 35 99 45 English
070 35 99 94 Dutch
070 35 98 66 French
Canada 1 712 432 2832 English
1 712 432 2844 French
Germany 01803 001 178 English
01570 248 8303 English (mobile)
01803 001 177 German
01570 248 8101 German (mobile)
Ireland 0818 270 007 English
Italy 848 390 166 English
848 391 819 Italian
Netherlands 0870 001 901 English
0870 001 909 Dutch
Poland 0801 003 543 English
0801 003 533 Polish
South Africa 087 550 0375 English
Spain 902 885 318 English
902 881 200 Spanish
Sweden 0939 2066 400 English
0939 2066 300 Swedish
Switzerland 0848 560 190 English
0848 560 378 French
0848 560 347 German
0848 560 385 Italian
United Kingdom (PowWowNow) 0844 4 73 73 73 English
United States 1 415 363 0833 English